# Benchmarking VLMs’ Reasoning About Persuasive Atypical Images

### v2

@@ Abstract
Vision language models (VLMs) have shown strong zero-shot generalization across various tasks, especially when integrated with large language models (LLMs). However, their ability to comprehend rhetorical and persuasive visual media, such as advertisements, remains understudied. Ads often employ atypical imagery, using surprising object juxtapositions to convey shared properties. For example, Fig. 1 (e) shows a beer with a feather-like texture. This requires advanced reasoning to deduce that this atypical representation signifies the beer’s lightness.
@@
Our findings reveal that: (1) VLMs lack advanced reasoning capabilities compared to LLMs; (2) simple, effective strategies can extract atypicality-aware information, leading to comprehensive image verbalization; (3) atypicality aids persuasive advertisement understanding. Code and data are available at github.com/sinamalakouti/PersuasiveAdVLMBenchmark
11footnotetext: These authors contributed equally to this work.
@@ 1 Introduction
Modern pretrained vision-language models (VLMs), such as LLaVA [27, 26], demonstrate strong visual understanding across various tasks such as recognition [25], and capabilities like zero-shot generalizability. However, there is a lack of in-depth study on VLMs’ ability to understand complex persuasive images, such as advertisements.
@@
Next, we investigate how atypicality influences understanding an ad’s message. We use the action-reason retrieval (ARR) task [18, 46], which requires models to identify the suggested action (e.g., “buy these chips") and its rationale (e.g., “because they are spicy"). However, to rigorously test the model’s reasoning, we introduce semantically challenging negative options, rather than mining hard negatives from other images [45, 19]. For example, we generate statements that include wrong action (e.g., “don’t buy these chips") or wrong rationale (e.g., “because they are sweet"). This prevents VLMs from ruling out negatives by merely comparing objects in the image and options. Our evaluation shows a significant performance drop when VLM is faced with hard negatives (e.g., LLaVA drops by 67.51%).
Finally, we hypothesize that deep atypicality understanding improves action-reason retrieval performance. We test this by proposing an atypicality-aware verbalization. We explore a few simple prompting strategies to construct a comprehensive, atypicality-sensitive ad verbalization, which is used to predict the atypicality statement. This statement, combined with the atypicality-aware verbalization, serves as input to an LLM classifier. This LLM integrates all the information to retrieve the final action-reason, effectively leveraging both visual understanding and high-level reasoning capabilities.
Our proposed framework achieves state-of-the-art performance on the ARR task. Interestingly, when a VLM is given both the image and our atypicality-aware verbalization, its performance on the ARR task actually declines (e.g., LLaVA($I+\mathcal{T}_{\mathcal{V}}$) shows a 1.71 point drop in performance compared to LLaVA($I$)) and it is significantly underperformed compared to LLM. This stark contrast highlights a critical gap: VLMs lack the advanced reasoning capabilities of LLMs when interpreting complex, atypical visual media. To summarize, our contributions are:
@@
We hope this work inspires the inclusion of persuasive ads in VLM benchmarks, foster development of robust models for complex visual media, and offer insights for creating more effective advertisements.
@@ 2 Related Works
Vision-language models. We benchmark pre-trained VLMs and LLMs on tasks involving atypicality and advertisement image understanding, focusing on their zero-shot reasoning capabilities. We also use pre-trained VLMs to verbalize advertisement images for LLMs. Given the substantial computational power required for training and fine-tuning large models, off-the-shelf, frozen models are typically used [39, 3]. Techniques to align visual and textual features without parameter updates include optimizing image encoders [39], inserting cross-attention layers [3], prompt learning [55, 54, 9, 21], and employing external transformers [32, 24, 13, 26]. However, direct application of models like [24, 33] may miss hidden messages by focusing more on visual context than semantics. Note that we restrict our experiments to the zero-shot inference setting.
@@
VLM evaluation. Various vision-language benchmarks have been proposed, including tasks like recognition [16], captioning [10], commonsense reasoning [40, 6], VCR [50], and compositional reasoning [37]. However, these benchmarks feature straightforward scenes that do not challenge the model’s reasoning abilities beyond literal imagery. Recently, ROME [53] generated counter-intuitive images focusing on 5 primitive common sense types (e.g., color and size) to challenge models’ ability in object, attribute, and spatial relation recognition. Similarly, WHOOPS! [7] addressed this by creating 500 synthetic scenes by placing ‘normal’ objects in unusual contexts (e.g., a snowplow in a desert and Einstein holding a smartphone). Atypical advertising images differ from those in WHOOPS! and ROME in two important ways: First, they are real ads created by human designers to intentionally convey a particular message (e.g., cigar replacing a bullet, to highlight the dangers of smoking), rather than simply to be unusual. This requires models to detect atypicality and atypical objects and reason about their impact on the ad’s meaning, making ad image understanding a more realistic and challenging benchmark for evaluating VLMs’ reasoning abilities. Second, atypical ads feature more categories of atypicality than placing an object out of context or altering primitive attributes, as WHOOPS! and ROME do, respectively.
@@ 3.1 Proposed Atypicality Understanding Tasks
Atypicality Statement Retrieval (ASR). ASR formulates atypicality inference as retrieving a statement describing relations between two objects. We represent atypicality $\mathcal{A}$ with templates $\mathcal{S_{\mathcal{A}}}$, defined in Fig. 3. A statement $s=(a,o^{p},o^{s})$ includes an atypicality type $a$, primary object $o^{p}$, and secondary object $o^{s}$ as described in Sec. 3.6 of [46]. In TR1 and TR2, $o^{p}$ is the object with a new texture, and $o^{s}$ is source of that texture. In OIO, $o^{p}$ is the object inside, and $o^{s}$ is the object outside. In OR, $o^{p}$ is the object placed in another’s context, and $o^{s}$ is the expected object.
@@
Atypical Object Recognition (AOR). To assess the VLMs’ fine-grained visual understanding, we introduce the task of recognizing the primary and secondary objects in an atypical image. Given an ad and the true atypicality, the model must generate the correct primary/secondary objects to complete the atypicality statement. AOR can be viewed as a fill-in-the-blank task (i.e., generative) based on the statement templates.
@@ 3.2 Proposed Approach
To explore the impact of atypicality in action-reason retrieval and compare the reasoning ability of VLMs and LLMs, we propose a three-steps in-context learning method : (i) Atypicality-aware image verbalization: Using LLaVA [27] and LLM, we generate a coherent verbalization $\mathcal{T}_{\mathcal{V}}$ sensitive to atypicality; (ii) Atypicality statement detection: We detect the atypicality statement $\hat{s}=(\hat{a},\hat{o}^{p},\hat{o}^{s})$; (iii) Action-reason retrieval: We combine steps (i) and (ii) to retrieve the final action-reason statement. Fig. 4 illustrates the method.
@@ 4 Experimental Setup
Baselines. We use LLaVA (‘vicuna-13b-v1.5’) [27], InstructBLIP (‘vicuna-13b-v0’) [13], MiniGPT4 [56] (‘vicuna-13b-v0’), and CLIP [33] (‘ViT-L/14@336px’ following [19]), and InternVL-Chat-V1-1 [11] as VLM baselines (and InternVL2-8B, LLaVA 1.6 in supp). LLaVA is our multimodal component due to its GPT-4-informed instruction tuning, state-of-the-art reasoning performance, and promptability [27]. We evaluate GPT-4V on a limited 250 examples, constrained by cost. We report BLIP-2 [24] (‘blip2-flan-t5-xl’) only for AOR as it failed to produce meaningful output when asked to return multiple options, i.e., MAC and multi-ARR (detail in supp).
Our analysis spans recent public LLMs, such as ‘vicuna-13b-v1.5’ (Vicuna) [38], ‘InternLM2-5-7b-chat’ (InternLM; see supp), and leading commercial models like GPT-3.5/4. We chose Vicuna as it is used in all VLMs (LLaVA, MiniGPT4, InstructBLIP) and InternLM as InternVL2-8B’s LM. We also compare GPT-4 and GPT-4V. We introduce CLIP ($I+T$) as a zero-shot baseline aligned with KAFA [19] but avoid direct comparison with KAFA as it is not publicly available. To assess our atypicality method, we compare against $V+T$ (verb. baseline), which includes basic image information (up to 5 objects, scene text).
Metrics. We use Precision, Recall, and F1-score to evaluate MAC (additional metrics in supp). For AOR, we assess sentence similarity between $s^{+}$ and $\hat{s}$ using ‘all-mpnet-base-v2’ [36], a state-of-the-art sentence embedding method. Common text-matching and accuracy metrics aren’t suitable for AOR since it is a generative task, and annotations can vary widely (e.g., ‘beer,’ ‘glass of beer,’ ‘beer glass’) due to human inconsistencies and typos. We report accuracy (denoted ‘Acc’) for single statement retrieval tasks, where the model returns only one statement per query (i.e., ASR and Single ARR). Top-k Acc and unranked Precision@k are the metrics for the multi-option ARR, with $Precision@k=\frac{min(k,\textit{\# of relevant statements in top k predictions})}{k}$. Note top-3 acc and prec@1 are the same.
@@
We validated our hard negatives by sampling 100 images and asking annotators to select correct statements (in supp).
Implementation. For GPT-4, GPT-3.5, and Vicuna temperature is set to 0. For LLaVA [26], BLIP-2 [24], InternVL models [11] and InternLM [8] we used the original setting. We applied 8-bit quantization for LLaVA, MiniGPT-4, Vicuna, InternVL and InternLM. All experiments were zero-shot and conducted on an NVIDIA 368 Quadro RTX A5000 GPU with 24 GB memory. Example prompts are in supp.
@@ 5.1 Atypicality Understanding Results
We assess opone-sour VLMs’ (LLaVA [26], InstructBLIP [13]) understanding of the atypicality and GPT-4V. Additionally, we evaluate the effectiveness of two prompting strategies for capturing atypicality: (1) $IN$ (Describe the image in detail.) and (2) $UH$ (What is unusual about the image?). These strategies are compared to the VLMs and the $V+T$ baseline across three LLMs (GPT-4, GPT-3.5, Vicuna). Table 1 summarizes the results on MAC and ASR, while AOR results are in Table 2.
@@ 5.2 Action-Reason Retrieval Results
We evaluate our proposed method using four LLMs (GPT-4, GPT-3.5, Vicuna, InternLM in supp) and compare them against VLMs (CLIP, LLaVA, InstructBLIP, InternVL-Chat-V1-1, GPT-4V, LLaVA1.6, and InternVL2-8B (in supp)). Table 3 presents the multi-option action-reason retrieval results (ARR) and their comparison with VLMs on the full-set. Table 4 exhibits ARR results on small-set when GPT-4V is the VLM in our pipeline. We adopted Vicuna as the public LLM because it’s used in LLaVA, InstructBLIP, and MiniGPT4. We use GPT-4 and GPT-4V as powerful LLM/VLM pairs.
@@
Table 3: ARR on Full-set . Predicted atypicality statement $\hat{s}$ uses the respective prompting (IN) and LLM (Vicuna, GPT-4), with all LLMs using LLaVA verbalization. $\mathcal{T}_{\mathcal{V}}$ combines $V+T+IN+UH$ using Vicuna or GPT-4. Best number per block and column is bolded. Precision@k is unranked. Task is Multi-ARR.
@@
Atypicality boosts persuasive visual media understanding. Table 3 demonstrates the effectiveness of our atypicality-aware verbalization method compared to VLMs and the $V+T$ verbalization baseline. While Vicuna($V+T$) lags behind LLaVA by 1.55 points on avg, indicating insufficient context from basic verbalization, our $\mathcal{T}_{\mathcal{V}}$ consistently improves $V+T$ performance. We observe gains of 3.8/2.8 and 1.93/2.46 on prec@2/prec@3 for Vicuna and GPT-4, respectively. Combining $\mathcal{T}_{\mathcal{V}}$ with $\hat{s}_{IN}$ yields the best results across all LLMs (GPT3.5 in supp), confirming the benefit of incorporating atypicality-aware verbalization and atypicality statement for understanding persuasive ads. Interestingly, Vicuna shows minimal gains when using GPT-4’s verbalization, suggesting that our strategy does not depend heavily on extensive LLMs like GPT-4.
@@ 5.3 Further Analysis & Ablation
Error analysis on ARR. In Fig. 5, VLMs (i.e., LLaVA and CLIP) do not make considerably more errors than GPT-4 ($\mathcal{T}_{\mathcal{V}}$) on ‘object swap’ since these negative options contain different objects from the ground-truth, making them easier for VLMs to identify. However, VLMs make substantially more errors than GPT-4($\mathcal{T}_{\mathcal{V}}$) on semantically incorrect negatives (i.e., ‘action alter,’ ‘reason alter,’ ‘statement alter,’ and ‘adjective alter’). This confirms VLMs mainly rely on visual elements (e.g., objects) rather than deeper reasoning (examples in supp).

### v3

@@ Abstract
Vision-language models (VLMs) have shown strong zero-shot generalization across various tasks, especially when integrated with large language models (LLMs). However, their ability to comprehend rhetorical and persuasive visual media, such as advertisements, remains understudied. Ads often employ atypical imagery, using surprising object juxtapositions to convey shared properties. For example, Fig. 1 (e) shows a beer with a feather-like texture. This requires advanced reasoning to deduce that this atypical representation signifies the beer’s lightness.
@@
Findings reveal that: (1) VLMs lack advanced reasoning capabilities compared to LLMs; (2) simple, effective strategies can extract atypicality-aware information, leading to comprehensive image verbalization; (3) atypicality aids persuasive ad understanding. Code and data is available at aysanaghazadeh.github.io/PersuasiveAdVLMBenchmark/
11footnotetext: Equal Contribution. Listing order is random.
@@ 1 Introduction
Modern pretrained vision-language models (VLMs) like LLaVA [29, 28] demonstrate strong visual understanding across various tasks such as recognition [26], and capabilities like zero-shot generalizability [37, 16]. However, there is a lack of in-depth study on VLMs’ ability to understand complex persuasive images such as advertisements.
@@
Next, we investigate how atypicality influences understanding an ad’s message. We use the action-reason retrieval (ARR) task [19, 48], which requires models to identify the suggested action (e.g., “buy these chips") and its rationale (e.g., “because they are spicy"). However, to rigorously test the model’s reasoning, we introduce semantically challenging negative options rather than mining hard negatives from other images [47, 20]. For example, we generate statements that include wrong action (e.g., “don’t buy these chips") or wrong rationale (e.g., “because they are sweet"). This prevents VLMs from ruling out negatives by merely comparing objects in the image and options. Our evaluation shows a significant performance drop when VLM is faced with hard negatives (e.g., LLaVA drops by 67.51%).
Finally, we hypothesize that deep atypicality understanding enhances action-reason retrieval performance. To test this, we propose an atypicality-aware verbalization. Using simple prompting strategies, we generate a comprehensive atypicality-sensitive ad verbalization to predict the corresponding atypicality statement. Then, an LLM integrates this statement with atypicality-aware verbalization to retrieve the final action-reason, effectively combining the benefit of both visual understanding and reasoning.
Our proposed framework achieves state-of-the-art performance on the ARR task. Interestingly, when a VLM is given both the image and our atypicality-aware verbalization, its performance on the ARR task declines (e.g., LLaVA($I+\mathcal{T}_{\mathcal{V}}$) shows a 1.71 point drop in performance compared to LLaVA($I$)) and it is significantly underperformed compared to LLM. This stark contrast highlights a critical gap: VLMs lack the advanced reasoning capabilities of LLMs when interpreting complex, atypical visual media. To summarize, our contributions are:
@@
We hope this work inspires the inclusion of persuasive ads in VLM benchmarks, fosters the development of robust models for complex visual media, and offers insights for creating more effective advertisements.
@@ 2 Related Works
Vision-language models. We benchmark pre-trained VLMs and LLMs on tasks involving atypicality and advertisement image understanding, focusing on their zero-shot reasoning capabilities. We use pre-trained VLMs to verbalize advertisement images for LLMs. Given the substantial computational power required for training and fine-tuning large models, off-the-shelf, frozen models are typically used [42, 3]. Techniques to align visual and textual features without parameter updates include optimizing image encoders [42], inserting cross-attention layers [3], prompt learning [57, 56, 9, 22], and employing external transformers [34, 25, 13, 28]. However, direct application of models like [25, 35] may miss hidden messages by focusing more on visual context than semantics. Note that we restrict our experiments to the zero-shot inference setting.
@@
VLM evaluation. Several vision-language benchmarks focus on tasks like recognition [17], captioning [10], commonsense reasoning [43, 6], VCR [52], and compositional reasoning [40]. However, these benchmarks typically involve simple scenes that fail to test the model’s reasoning abilities beyond the basic interpretation of literal imagery. Recently, ROME [55] generated counter-intuitive images focusing on 5 primitive common sense types (e.g., color and size) to challenge models’ ability in object, attribute, and spatial relation recognition. Similarly, WHOOPS! [7] addressed this by creating 500 synthetic scenes by placing ‘normal’ objects in unusual contexts (e.g., a snowplow in a desert and Einstein holding a smartphone). Atypical advertising images differ from those in WHOOPS! and ROME in two important ways: First, they are real ads created by human designers to intentionally convey a particular message (e.g., cigar replacing a bullet, to highlight the dangers of smoking), rather than merely aiming to be unusual. This requires models to detect atypicality and atypical objects and reason about their impact on the ad’s message. Ad image understanding becomes a more realistic and challenging benchmark for evaluating VLMs’ reasoning ability. Second, atypical ads involve more types of atypicality than simply placing an object out of context, as in WHOOPS!, or altering primitive attributes, as in ROME.
@@ 3.1 Proposed Atypicality Understanding Tasks
Atypicality Statement Retrieval (ASR). ASR frames atypicality inference as retrieving a statement describing relations between two objects. Atypicality $\mathcal{A}$ is presented using templates $\mathcal{S_{\mathcal{A}}}$, as defined in Fig. 3. Each statement $s=(ao^{p},o^{s})$ includes an atypicality type $a$, primary object $o^{p}$, and secondary object $o^{s}$ as described in Sec. 3.6 of [48]. For TR1 and TR2, $o^{p}$ is the object with the new texture, and $o^{s}$ is the texture source. For OIO, $o^{p}$ is the object inside, and $o^{s}$ is the object outside. For OR, $o^{p}$ is the object placed in the context of another, and $o^{s}$ is the expected object.
@@
Atypical Object Recognition (AOR). To assess the VLMs’ fine-grained visual understanding, we propose recognizing the primary and secondary objects in an atypical image. Given ad image and true atypicality, the goal is to generate the correct primary/secondary objects to complete the atypicality statement. AOR functions as a fill-in-the-blank task (i.e., generative) using using statement templates.
@@ 3.2 Proposed Approach
To explore the impact of atypicality in action-reason retrieval and compare the reasoning abilities of VLMs and LLMs, we propose an in-context learning method consisting of three steps: (i) Atypicality-aware image verbalization: Using LLaVA [29] and LLM, we generate a coherent verbalization $\mathcal{T}_{\mathcal{V}}$ sensitive to atypicality; (ii) Atypicality statement detection: We detect the atypicality statement $\hat{s}=(\hat{a},\hat{o}^{p},\hat{o}^{s})$; (iii) Action-reason retrieval: We retrieve the final action-reason statement using (i) and (ii). Fig. 4 illustrates the proposed method.
@@ 4 Experimental Setup
Baselines. We use LLaVA (‘vicuna-13b-v1.5’) [29], InstructBLIP (‘vicuna-13b-v0’) [13], MiniGPT4 [58] (‘vicuna-13b-v0’), and CLIP [35] (‘ViT-L/14@336px’ following [20]), and ‘InternVL-Chat-V1-1’ (denoted as InternVL-V1) [11] as VLM baselines (and ‘InternVL2-8B’, LLaVA 1.6 in supp Tab. 4). LLaVA is our multimodal component due to its GPT-4-informed instruction tuning, state-of-the-art reasoning, and promptability [29]. We evaluate GPT-4V on a limited 250 examples, constrained by cost. We report BLIP-2 [25] (‘blip2-flan-t5-xl’) only for AOR as it failed to produce meaningful output for multi-option tasks (i.e., MAC and multi-ARR; detail in supp sec. 2.2).
Our analysis spans recent public LLMs, such as ‘vicuna-13b-v1.5’ (Vicuna) [41], ‘InternLM2-5-7b-chat’ (InternLM; see supp sec. 2.2), and leading commercial models like GPT-3.5/4. We chose Vicuna as it is used in all VLMs (LLaVA, MiniGPT4, InstructBLIP) and InternLM as InternVL2-8B’s LM. We also compare GPT-4 and GPT-4V. We introduce CLIP ($I+T$) as a zero-shot baseline aligned with KAFA [20] but avoid direct comparison with KAFA as it is not publicly available. To assess our atypicality method, we compare against $V+T$ (verb. baseline), which includes basic image information (up to 5 objects, scene text).
Metrics. We use Precision, Recall, and F1-score to evaluate MAC (additional metrics in supp Tab. 1). For AOR, we assess sentence similarity between $s^{+}$ and $\hat{s}$ using ‘all-mpnet-base-v2’ [39], a state-of-the-art sentence embedding method. Common text-matching and accuracy metrics aren’t suitable for AOR since it is a generative task, and annotations can vary widely (e.g., ‘beer,’ ‘glass of beer,’ ‘beer glass’) due to human inconsistencies and typos. We report accuracy (denoted ‘Acc’) for single statement retrieval tasks, where the model returns only one statement per query (i.e., ASR and Single ARR). Top-k Acc and unranked Precision@k are the metrics for the multi-option ARR, with $Precision@k=\frac{min(k,\textit{\# of relevant statements in top k predictions})}{k}$. Note top-3 acc and prec@1 are the same.
@@
We validated our hard negatives by sampling 100 images and asked human annotators to select all correct statements (results in supp sec. 2.2).
Implementation. For GPT-4, GPT-3.5, and Vicuna temperature is set to 0. For LLaVA [28], BLIP-2 [25], InternVL models [11] and InternLM [8] we used the original setting. We applied 8-bit quantization for LLaVA, MiniGPT-4, Vicuna, InternVL and InternLM. All experiments were zero-shot and conducted on an NVIDIA Quadro RTX A5000. Example prompts are in supp.
@@ 5.1 Atypicality Understanding Results
We assess open-source VLMs’ (LLaVA [28], InstructBLIP [13]) understanding of the atypicality and GPT-4V. Additionally, we evaluate the effectiveness of two prompting strategies for capturing atypicality: (1) $IN$ (Describe the image in detail.) and (2) $UH$ (What is unusual about the image?). These strategies are compared to the VLMs and the $V+T$ baseline across three LLMs (GPT-4, GPT-3.5, Vicuna). Table 1 summarizes the results on MAC and ASR, while AOR results are in Table 2.
@@ 5.2 Action-Reason Retrieval Results
We evaluate our proposed method using three LLMs (GPT-4, GPT-3.5, Vicuna, InternLM in supp) and compare them against VLMs (CLIP, LLaVA, InstructBLIP, InternVL-V1, GPT-4V, and LLaVA1.6 and InternVL2-8B in supp.). Table 3 presents the multi-option action-reason retrieval results (ARR) and their comparison against VLMs on the full-set. Table 4 demonstrates ARR results on small-set when GPT-4V is the VLM in our pipeline. We adopted Vicuna as the public LLM as it is used in all VLMs (LLaVA, InstructBLIP, and MiniGPT4). We use GPT-4 and GPT-4V as powerful LLM/VLM pairs.
@@
Table 3: ARR on Full-set . Predicted atypicality statement $\hat{s}$ uses the respective prompting (IN) and LLM (Vicuna, GPT-4), with all LLMs using LLaVA verbalization. $\mathcal{T}_{\mathcal{V}}$ combines $V+T+IN+UH$ using Vicuna or GPT-4. Best number per block and column is bolded. Task is Multi-ARR. Precision@k is unranked.
@@
Atypicality boosts persuasive visual media understanding. Table 3 highlights the effectiveness of our atypicality-aware verbalization compared to VLMs and the basic $V+T$ baseline. Vicuna($V+T$) falls 1.55 points behind LLaVA on avg, reflecting insufficient context from basic verbalization. In contrast, $\mathcal{T}_{\mathcal{V}}$ consistently improves $V+T$, with gains of 3.8/2.8 (prec@2/prec@3) for Vicuna and 1.93/2.46 for GPT-4. Combining $\mathcal{T}_{\mathcal{V}}$ with $\hat{s}_{IN}$ achieves the best results across all LLMs (GPT3.5 in supp), confirming the benefit of incorporating atypicality-aware verbalization and atypicality. Notably, Vicuna shows minimal improvement with GPT-4’s verbalization, indicating our strategy does not strictly rely on extensive LLMs like GPT-4.
@@ 5.3 Further Analysis & Ablation
Error analysis on ARR. In Fig. 5, VLMs (i.e., LLaVA and CLIP) perform comparably to ($\mathcal{T}_{\mathcal{V}}$) on ‘object swap’ as these negative options include different objects from the ground-truth, making them easy for VLMs to identify. However, VLMs make substantially more errors than GPT-4($\mathcal{T}_{\mathcal{V}}$) on semantically incorrect negatives (i.e., ‘action alter,’ ‘reason alter,’ ‘statement alter,’ and ‘adjective alter’ ). This confirms VLMs mainly rely on visual elements (e.g., objects) rather than deeper reasoning (examples in supp).